# Cost-aware LLM-based Online Dataset Annotation

### v1

@@ 4.1 Experiments on the MMLU Dataset
We first evaluate CaMVo on the MMLU dataset Hendrycks et al. (2021a, b), a challenging multiple-choice benchmark spanning 57 diverse subjects including mathematics, U.S. history, law, and computer science. MMLU is well-suited to our setting, as it demands broad world knowledge and strong reasoning capabilities—conditions under which majority voting is particularly effective. To reduce computational cost, we restrict our evaluation to the test split, which contains 14,042 instances.
@@
We use the following LLMs: Claude 3 Sonnet and Haiku from Anthropic Anthropic (2024), GPT-4o, o3-mini, and o1-mini from OpenAI OpenAI (2024), and LLaMA-3.3 and LLaMA-3.1 from Meta Meta (2024). All models are queried using temperature 0.35 and top-$p=1$, where applicable. To extract contextual embeddings for CaMVo, we use the 384-dimensional sentence transformer all-MiniLM-L6-v2 Wang et al. (2020). For computational efficiency, we approximate the confidence $\delta_{\mathcal{A}}(\bm{L},\bm{\omega})$ using the cumulative distribution function of a Beta distribution. Further implementation details and experimental setup are provided in Appendix D.
@@ Results.
Table 2 (Left) reports the accuracy and cost of individual LLMs, as well as two baselines: Majority Vote, which aggregates all LLMs using weights proportional to their true accuracy, and the Baseline Method, which corresponds to Algorithm 2. Cost corresponds to the average input token cost when labeling the dataset and is reported in dollars per million input tokens. Accuracy reflects the percentage of data instances correctly labeled. CaMVo’s performance under varying confidence thresholds $\delta$ and minimum vote counts $k_{\min}\in\{1,3\}$ is shown in Table 3. Note that we include $k_{\min}=3$ in our experiments as model parameters do not get updated when only a single LLM is queried, a scenario that can occur under $k_{\min}=1$. As a result, users may prefer to avoid setting $k_{\min}=1$. Moreover, for $k_{\min}=2$, the selected pair of LLMs will always yield a majority vote in favor of the LLM with the higher weight, limiting the informativeness of the voting outcome. The algorithm is configured with $\alpha=0.25$, $\lambda_{R}=1$, and $\lambda_{L}=1$. The Target Accuracy column reflects the minimum accuracy CaMVo must exceed to satisfy the threshold $\delta$, and is computed as $\delta\times\text{(Majority Vote Accuracy)}=\delta\times 88.18\%$.
@@
CaMVo $\delta$ Target Acc. (%) Acc. (%) $k_{\min}=1$ Cost $k_{\min}=1$ Acc. (%) $k_{\min}=3$ Cost $k_{\min}=3$
@@
Figure 1: (Left) Cost–accuracy trade-off for MMLU dataset: gray dots show every LLM subset via weighted majority voting, yellow dots trace their Pareto-optimal frontier, blue markers are CaMVo at $k_{\min}=1$, green markers at $k_{\min}=3$, and the red marker denotes the Baseline Method. (Right) Cumulative average accuracy and cost of CaMVo with $\delta=0.96$, $k_{\min}=1$ over rounds.
Figure 1 (Left) illustrates the cost–accuracy trade-off of All Subsets versus CaMVo. Each gray point represents one of the $12^{K}\!-\!$ 1 possible LLM subsets voted via ground-truth accuracies, while the yellow curve depicts those that are Pareto-optimal. CaMVo’s results appear as blue markers for $k_{\min}=1$ and green markers for $k_{\min}=3$. The red marker denotes the Baseline Method. Remarkably, even without any a priori knowledge of LLM performance, or any pre-training, CaMVo consistently tracks, and sometimes surpasses the Pareto frontier, demonstrating its ability to approximate optimal cost–accuracy trade-offs in an online manner.
Figure 1 (Right) shows CaMVo’s cumulative average accuracy (blue) and cost (red) for $\delta=0.96$, $k_{\min}=1$; the green horizontal line indicates the target accuracy. In early rounds, CaMVo explores larger, more expensive subsets, yielding both high cost and high accuracy. As the LCB estimates converge, the algorithm rapidly shifts to smaller, cheaper subsets that still satisfy the accuracy threshold. Cost declines steeply while accuracy stabilizes just above the target, illustrating CaMVo’s ability to quickly identify and exploit the most cost-effective ensembles without sacrificing labeling quality. Additional plots with different parameters are provided in Appendix D.
@@ 4.2 Experiments on the IMDB Movie Reviews Dataset
We next test CaMVo on the IMDB Movie Reviews dataset Maas et al. (2011), a balanced binary-sentiment benchmark of 50,000 movie reviews. As before, we compare CaMVo against each individual LLM, a full-ensemble Majority Vote, and the Baseline Method (Algorithm 2).
@@
We employ Anthropic’s Claude 3-5 Haiku Anthropic (2024); OpenAI’s GPT-4o, o3-mini, GPT-4o-mini, and o1-mini OpenAI (2024); and Meta’s LLaMA-3.3 and LLaMA-3.1 Meta (2024). All queries use temperature = 0.25 and top-$p=1$, where applicable. We extract 384-dimensional contextual embeddings with all-MiniLM-L6-v2 Wang et al. (2020) and approximate the confidence bound $\delta_{\mathcal{A}}(\bm{L},\bm{\omega})$ via the Beta-CDF, as in §4.1.
@@ Results.
Table 2 (Right) reports the accuracy and cost (in dollars per million input tokens) of each LLM and the two baselines. The baseline underperforms the best individual model (95.68% vs. 95.61%) despite incurring a significantly higher cost. This is partly due to the relative ease of the IMDB Movie Reviews dataset, where individual LLMs already achieve high accuracy, limiting the marginal benefit of ensembling. As noted by Li et al. (2024b), ensemble gains are most pronounced on harder tasks. Additionally, Trad and Chehab (2024) highlight that large performance gaps among models can reduce ensemble effectiveness, making smaller, selective subsets preferable in such cases.
Table 4 presents CaMVo’s accuracy–cost trade-off across various thresholds $\delta$ and $k_{\min}\in\{1,3\}$. CaMVo’s hyperparameters are $\alpha=0.7$, $\lambda_{R}=5$, and $\lambda_{L}=1$; and the Target Accuracy is computed similarly as $\delta\times 95.62\%$. Across all configurations, CaMVo meets or exceeds its target accuracy. Further, CaMVo achieves less than half the cost (when $\delta=0.997$ and $k_{\min}=1$) at a slightly lower accuracy of $95.45\%$ compared to the baseline, confirming its practicality for large-scale sentiment annotation without any pre-training or ground-truth labels.
CaMVo $\delta$ Target Acc. (%) Acc. (%) $k_{\min}=1$ Cost $k_{\min}=1$ Acc. (%) $k_{\min}=3$ Cost $k_{\min}=3$
@@
Figure 2: (Left) Cost–accuracy trade-off for IMDB dataset: gray dots show every LLM subset via weighted majority voting, yellow dots trace their Pareto-optimal frontier, blue markers are CaMVo at $k_{\min}=1$, green markers at $k_{\min}=3$, and the red marker denotes the Baseline Method. (Right) Empirical average accuracy and cost of CaMVo with $\delta=0.995$, $k_{\min}=1$ over rounds.
Figure 2 (Left) presents the analogous comparison of Figure 1 (Left) on the IMDB sentiment task. As before, gray points and the yellow Pareto-frontier points show all possible subset combinations, while blue and green markers plot CaMVo at $k_{\min}=1$ and $3$, respectively. The red marker denotes the Baseline Method. CaMVo closely matches the Pareto front in the low-cost regime (cost $<1$), but lags behind in higher-cost regions. This exposes a key limitation: when majority voting with additional LLMs is ineffective, CaMVo’s reliance on the independence assumption, which suggests that aggregating more LLMs improves accuracy; can lead to suboptimal performance.
Figure 2 (Right) plots CaMVo’s cumulative average accuracy (blue) and cost (red) on IMDB with $\delta=0.995$ and $k_{\min}=1$; the green line marks the target accuracy. As before, early rounds involve querying larger, costlier ensembles to robustly explore each model’s performance. Once the lower-confidence bounds stabilize, CaMVo swiftly transitions to minimal-cost subsets that still meet the accuracy requirement. This demonstrates CaMVo’s rapid convergence to cost-effective model combinations without compromising annotation quality. Additional results for other parameter settings appear in Appendix E.
@@ 5 Limitations
Our work relies on the assumption that the outputs of LLMs are independent of each other. Under this assumption, aggregating any subset of models with individual accuracy above $50\%$ strictly improves majority-vote performance. In practice; i.e., on the IMDB sentiment task (§4.2), LLM outputs can be highly correlated, and majority voting may underperform the best single model. Consequently, CaMVo inherits these failures and can yield lower ensemble accuracy when independence is violated. Nevertheless, even in such regimes CaMVo still achieves the user-specified accuracy threshold while reducing cost relative to the full-ensemble baseline. This is mostly due to the fact that our results are relative to the full-ensemble baseline which also suffers from the same issue. Extending CaMVo by accounting for inter-model correlations (e.g., via joint confidence estimation or diversity-aware selection) to better find the optimal subset is a promising and challenging direction for future work.
@@ 6 Conclusions
We have introduced Cost-aware Majority Voting (CaMVo), the first fully online framework for LLM-based dataset labeling that jointly adapts both vote weights and the subset of models queried on a per-instance basis. By combining a LinUCB-style contextual bandit with a Bayesian Beta-mixture confidence estimator, CaMVo estimates a lower bound on each LLM’s correctness probability for the given input and selects the minimal-cost ensemble that meets a user-specified accuracy threshold.
Empirical results on the MMLU and IMDB benchmarks demonstrate that CaMVo matches or exceeds full-ensemble majority-vote accuracy while reducing labeling cost. On MMLU, CaMVo even surpasses the true Pareto frontier of all possible weighted subsets—despite having no prior knowledge of individual model performance. These findings establish CaMVo as a practical solution for cost-efficient, automated annotation in dynamic labeling environments without any ground-truth labels or offline training.
Our analysis assumes independence among LLM outputs, which can be violated in practice and may degrade ensemble gains. Nonetheless, CaMVo still enforces the user’s accuracy target and delivers significant cost savings even under these conditions. Future work will explore diversity-aware selection and joint confidence models to mitigate correlated errors. We will also extend CaMVo to support iterative relabeling, allowing previously annotated instances to be revisited and refined as additional contextual information becomes available.
Acknowledgments
This work was supported in part by the National Science Foundation through Grant CCF-2007834, by the Office of Naval Research through Grant #N00014-23-1-2275, and by the CyLab Enterprise Security Initiative. The work of C. Tekin was supported by TUBITAK 2024 Incentive Award and TUBA-GEBIP 2023 Award.
@@ Acknowledgments
Proof.
Assume that each LLM $l_{i}\in\mathcal{A}$ produces a correct label with probability at least $L_{i}$, independently of other models. Define the random variable $Z_{i}\sim\text{Bernoulli}(L_{i})$ to represent whether model $l_{i}$ correctly labels the data instance, where $\mathbb{E}[Z_{i}]\geq L_{i}$. The total weight of LLMs that output the correct label is given by:
@@ Appendix B Estimating the Shape Parameters of the Beta Distribution
$\bar{q}_{i,1}$ $=\frac{1}{|S_{i,t}|}\sum_{s\in S_{i,t}}q_{i}(e_{s},s),\quad$ $v_{i,1}^{2}=\frac{1}{|S_{i,t}|}\sum_{s\in S_{i,t}}(q_{i}(e_{s},s)-\bar{q}_{i,1})^{2}$ (14)
$\bar{q}_{i,0}$ $=\frac{1}{t-|S_{i,t}|}\sum_{s\in[t]\setminus S_{i,t}}q_{i}(e_{s},s),\quad$ $v_{i,0}^{2}=\frac{1}{t-|S_{i,t}|}\sum_{s\in[t]\setminus S_{i,t}}(q_{i}(e_{s},s)-\bar{q}_{i,0})^{2}$ (15)

### v2

@@ 4.1 Experiments on the MMLU Dataset
We first evaluate CaMVo on the MMLU dataset hendrycks2021ethics; hendryckstest2021, a challenging multiple-choice benchmark spanning 57 diverse subjects including mathematics, U.S. history, law, and computer science. MMLU is well-suited to our setting, as it demands broad world knowledge and strong reasoning capabilities—conditions under which majority voting is particularly effective. To reduce computational cost, we restrict our evaluation to the test split, which contains 14,042 instances.
@@
We use the following LLMs: Claude 3 Sonnet and Haiku from Anthropic claude3sonnet, GPT-4o, o3-mini, and o1-mini from OpenAI openai2024models, and LLaMA-3.3 and LLaMA-3.1 from Meta llama2024models. All models are queried using temperature 0.35 and top-$p=1$, where applicable. To extract contextual embeddings for CaMVo, we use the 384-dimensional sentence transformer all-MiniLM-L6-v2 wang2020minilm. For computational efficiency, we approximate the confidence $\delta_{\mathcal{A}}(\bm{L},\bm{\omega})$ using the cumulative distribution function of a Beta distribution. Further implementation details and experimental setup are provided in Appendix D.
@@ Results.
Table 2 (Left) reports the accuracy and cost of individual LLMs, as well as two baselines: Majority Vote, which aggregates all LLMs using weights proportional to their true accuracy, and the Baseline Method, which corresponds to Algorithm 2. Cost corresponds to the average input token cost when labeling the dataset and is reported in dollars per million input tokens. Accuracy reflects the percentage of data instances correctly labeled. CaMVo’s performance under varying confidence thresholds $\delta$ and minimum vote counts $k_{\min}\in\{1,3\}$ is shown in Table 3. Note that we include $k_{\min}=3$ in our experiments as model parameters do not get updated when only a single LLM is queried, a scenario that can occur under $k_{\min}=1$. As a result, users may prefer to avoid setting $k_{\min}=1$. Moreover, for $k_{\min}=2$, the selected pair of LLMs will always yield a majority vote in favor of the LLM with the higher weight, limiting the informativeness of the voting outcome. The algorithm is configured with $\alpha=0.25$, $\lambda_{R}=1$, and $\lambda_{L}=1$. The Target Accuracy column reflects the targeted relative accuracy, which is the minimum relative accuracy CaMVo must exceed to satisfy the threshold $\delta$, and is computed as $\delta\times\text{(Majority Vote Accuracy)}=\delta\times 88.18\%$.
@@
CaMVo $\delta$
Target
Acc. (%)
Acc. (%)
$k_{\min}=1$
Cost
$k_{\min}=1$
Acc. (%)
$k_{\min}=3$
Cost
$k_{\min}=3$
@@
Figure 1: (Left) Cost–accuracy trade‐off for MMLU dataset: gray dots show every LLM subset via weighted majority voting, yellow dots trace their Pareto‐optimal frontier, blue markers are CaMVo at $k_{\min}=1$, green markers at $k_{\min}=3$, cyan markers denote the individual single LLMs, and the red marker denotes the Baseline Method. (Right) Cumulative average accuracy and cost of CaMVo with $\delta=0.96$, $k_{\min}=1$ over rounds.
Figure 1 (Left) illustrates the cost–accuracy trade-off of All Subsets versus CaMVo. Each gray point represents one of the $2^{K}\!-\!1$ possible LLM subsets voted via ground‐truth accuracies, while the yellow curve depicts those that are Pareto‐optimal. CaMVo’s results appear as blue markers for $k_{\min}=1$ and green markers for $k_{\min}=3$. The red marker denotes the Baseline Method. Remarkably, even without any a priori knowledge of LLM performance, or any pre-training, CaMVo consistently tracks, and sometimes surpasses the Pareto frontier, demonstrating its ability to approximate optimal cost–accuracy trade‐offs in an online manner.
Figure 1 (Right) shows CaMVo’s cumulative average accuracy (blue) and cost (red) for $\delta=0.96$, $k_{\min}=1$; the green horizontal line indicates the target accuracy. In early rounds, CaMVo explores larger, more expensive subsets, yielding both high cost and high accuracy. As the LCB estimates converge, the algorithm rapidly shifts to smaller, cheaper subsets that still satisfy the accuracy threshold. Cost declines steeply while accuracy stabilizes just above the target, illustrating CaMVo’s ability to quickly identify and exploit the most cost‐effective ensembles without sacrificing labeling quality. Additional plots with different parameters are provided in Appendix D.
@@ 4.2 Experiments on the IMDB Movie Reviews Dataset
We next test CaMVo on the IMDB Movie Reviews dataset maas-EtAl:2011:ACL-HLT2011, a balanced binary‐sentiment benchmark of 50,000 movie reviews. As before, we compare CaMVo against each individual LLM, a full‐ensemble Majority Vote, and the Baseline Method (Algorithm 2).
@@
We employ Anthropic’s Claude 3-5 Haiku claude3sonnet; OpenAI’s GPT-4o, o3-mini, GPT-4o-mini, and o1-mini openai2024models; and Meta’s LLaMA-3.3 and LLaMA-3.1 llama2024models. All queries use temperature = 0.25 and top-$p=1$, where applicable. We extract 384-dimensional contextual embeddings with all-MiniLM-L6-v2 wang2020minilm and approximate the confidence bound $\delta_{\mathcal{A}}(\bm{L},\bm{\omega})$ via the Beta‐CDF, as in §4.1.
@@ Results.
Table 2 (Right) reports the accuracy and cost (in dollars per million input tokens) of each LLM and the two baselines. The baseline underperforms the best individual model (95.68% vs. 95.61%) despite incurring a significantly higher cost. This is partly due to the relative ease of the IMDB Movie Reviews dataset, where individual LLMs already achieve high accuracy, limiting the marginal benefit of ensembling. As noted by li2024more, ensemble gains are most pronounced on harder tasks. Additionally, trad2024ensemble highlight that large performance gaps among models can reduce ensemble effectiveness, making smaller, selective subsets preferable in such cases.
Table 4 presents CaMVo’s accuracy–cost trade‐off across various thresholds $\delta$ and $k_{\min}\in\{1,3\}$. CaMVo’s hyperparameters are $\alpha=0.7$, $\lambda_{R}=5$, and $\lambda_{L}=1$; and the Target Accuracy is computed similarly as $\delta\times 95.62\%$. Across all configurations, CaMVo meets or exceeds its target accuracy. Further, CaMVo achieves less than half the cost (when $\delta=0.997$ and $k_{\min}=1$) at a slightly lower accuracy of $95.45\%$ compared to the baseline, confirming its practicality for large‐scale sentiment annotation without any pre‐training or ground-truth labels.
CaMVo $\delta$
Target
Acc. (%)
Acc. (%)
$k_{\min}=1$
Cost
$k_{\min}=1$
Acc. (%)
$k_{\min}=3$
Cost
$k_{\min}=3$
@@
Figure 2: (Left) Cost–accuracy trade‐off for IMDB dataset: gray dots show every LLM subset via weighted majority voting, yellow dots trace their Pareto‐optimal frontier, blue markers are CaMVo at $k_{\min}=1$, green markers at $k_{\min}=3$, cyan markers denote the individual single LLMs, and the red marker denotes the Baseline Method. (Right) Empirical average accuracy and cost of CaMVo with $\delta=0.995$, $k_{\min}=1$ over rounds.
Figure 2 (Left) presents the analogous comparison of Figure 1 (Left) on the IMDB sentiment task. As before, gray points and the yellow Pareto-frontier points show all possible subset combinations, while blue and green markers plot CaMVo at $k_{\min}=1$ and $3$, respectively. The red marker denotes the Baseline Method. CaMVo closely matches the Pareto front in the low-cost regime (cost <1<1), but lags behind in higher-cost regions. This exposes a key limitation: when majority voting with additional LLMs is ineffective, CaMVo’s reliance on the independence assumption, which suggests that aggregating more LLMs improves accuracy; can lead to suboptimal performance.
Figure 2 (Right) plots CaMVo’s cumulative average accuracy (blue) and cost (red) on IMDB with $\delta=0.995$ and $k_{\min}=1$; the green line marks the target accuracy. As before, early rounds involve querying larger, costlier ensembles to robustly explore each model’s performance. Once the lower‐confidence bounds stabilize, CaMVo swiftly transitions to minimal‐cost subsets that still meet the accuracy requirement. This demonstrates CaMVo’s rapid convergence to cost‐effective model combinations without compromising annotation quality. Additional results for other parameter settings appear in Appendix E.
4.3 Additional Experiments
We present additional experimental results on the AG News Classification Dataset in Appendix F. Appendix G introduces CCaMVo (Correlated CaMVo), a practical extension of CaMVo that estimates subset confidence by modeling correlations among LLMs, thereby relaxing the independence assumption. We provide the corresponding algorithm, experiments, and a comparison with CaMVo, showing that accounting for correlations yields only marginal improvements in accuracy. Finally, Appendix H reports a sensitivity analysis evaluating how CaMVo’s performance degrades under varying levels of correlation among LLM predictions.
@@ 5 Limitations
Our work relies on the assumption that the outputs of LLMs are independent of each other. Under this assumption, aggregating any subset of models with individual accuracy above $50\%$ strictly improves majority‐vote performance. In practice; i.e., on the IMDB sentiment task (§4.2), LLM outputs can be highly correlated, and majority voting may underperform the best single model. Consequently, CaMVo inherits these failures and can yield lower ensemble accuracy when independence is violated. Nevertheless, even in such regimes CaMVo still achieves the user‐specified accuracy threshold while reducing cost relative to the full‐ensemble baseline. This is mostly due to the fact that our results are relative to the full‐ensemble baseline which also suffers from the same issue.
Appendix G introduces CCaMVo (Correlated CaMVo), a practical extension that estimates subset confidence thorugh estimating the correlation matrix among LLMs; this approach yields only marginal gains in accuracy. More broadly, extending CaMVo to account for inter-model correlations in a principled manner—e.g., via joint confidence estimation or diversity-aware subset selection—represents a promising and challenging direction for future work.
@@ 6 Conclusions
We have introduced Cost‐aware Majority Voting (CaMVo), the first fully online framework for LLM‐based dataset labeling that jointly adapts both vote weights and the subset of models queried on a per‐instance basis. By combining a LinUCB‐style contextual bandit with a Bayesian Beta‐mixture confidence estimator, CaMVo estimates a lower bound on each LLM’s correctness probability for the given input and selects the minimal‐cost ensemble that meets a user‐specified accuracy threshold.
Empirical results on the MMLU and IMDB benchmarks demonstrate that CaMVo matches or exceeds full‐ensemble majority‐vote accuracy while reducing labeling cost. On MMLU, CaMVo even surpasses the true Pareto frontier of all possible weighted subsets—despite having no prior knowledge of individual model performance. These findings establish CaMVo as a practical solution for cost‐efficient, automated annotation in dynamic labeling environments without any ground‐truth labels or offline training.
Our analysis assumes independence among LLM outputs, which can be violated in practice and may degrade ensemble gains. Nonetheless, CaMVo still enforces the user’s accuracy target and delivers significant cost savings even under these conditions.
Our method can be naturally extended beyond classification tasks in several ways. For regression, LLM outputs can be aggregated via a weighted average or median, with weights updated based on how closely each LLM’s output aligns with the aggregate. For ranking tasks, each LLM can assign scores to items, and a weighted combination of these scores would produce a final ranking; weights can then be adjusted according to agreement of the LLM’s ranking generated from its scores with the aggregated ranking. These extensions involve only minor modifications to the aggregation and weight update steps, making them straightforward to integrate into our existing framework. Future work will explore diversity‐aware selection and joint confidence models to mitigate correlated errors. We will also extend CaMVo to support iterative relabeling, allowing previously annotated instances to be revisited and refined as additional contextual information becomes available.
7 Acknowledgements
We thank Dr. Samarth Gupta for his thoughtful insights and stimulating discussions that helped shape the initial direction of this paper. This work was supported in part by the Office of Naval Research Grant # N00014-23-1-2275 and by CyLab Enterprise Security Initiative. The work of Cem Tekin was supported in part by the Scientific and Technological Research Council of Türkiye (TÜBİTAK) under Grant 124E065; by the Turkish Academy of Sciences Distinguished Young Scientist Award Program (TÜBA-GEBİP-2023); and by TÜBİTAK 2024 Incentive Award.
@@ 7 Acknowledgements
In order to find a lower bound on the probability of correct labeling, we consider the worst-case where each LLM’s probability of correct labeling is exactly equal its lower bound. Hence, assume that each LLM $l_{i}\in\mathcal{A}$ produces a correct label with probability $L_{i}$, independently of other models. Define the random variable $Z_{i}\sim\text{Bernoulli}(L_{i})$ to represent whether model $l_{i}$ correctly labels the data instance, where $\mathbb{E}[Z_{i}]\geq L_{i}$. The total weight of LLMs that output the correct label is given by:
@@ Appendix B Estimating the Shape Parameters of the Beta Distribution
$\bar{q}_{i,1}$ $=\frac{1}{|S_{i,t}|}\sum_{s\in S_{i,t}}q_{i,s}(\bf{e}_{s}),\qquad$ $v_{i,1}^{2}=\frac{1}{|S_{i,t}|}\sum_{s\in S_{i,t}}(q_{i,s}(\bm{e}_{s})-\bar{q}_{i,1})^{2}$ (14)
$\bar{q}_{i,0}$ $=\frac{1}{t-|S_{i,t}|}\sum_{s\in[t]\setminus S_{i,t}}q_{i,s}(\bm{e}_{s}),\quad$ $v_{i,0}^{2}=\frac{1}{t-|S_{i,t}|}\sum_{s\in[t]\setminus S_{i,t}}(q_{i,s}(\bm{e}_{s})-\bar{q}_{i,0})^{2}$ (15)